# Direction-Dependent Conduction Polarity in Altermagnetic CrSb

### v1

@@
Banik Rai Department of Condensed Matter and Materials Physics, S. N. Bose National Centre for Basic Sciences, Salt Lake City, Kolkata-700106, India These authors contributed equally to this work. Krishnendu Patra Department of Condensed Matter and Materials Physics, S. N. Bose National Centre for Basic Sciences, Salt Lake City, Kolkata-700106, India These authors contributed equally to this work. Satyabrata Bera School of Physical Sciences, Indian Association for the Cultivation of Science, Jadavpur, Kolkata-700032, India Kakan Deb Department of Condensed Matter and Materials Physics, S. N. Bose National Centre for Basic Sciences, Salt Lake City, Kolkata-700106, India Mintu Mondal School of Physical Sciences, Indian Association for the Cultivation of Science, Jadavpur, Kolkata-700032, India Priya Mahadevan Department of Condensed Matter and Materials Physics, S. N. Bose National Centre for Basic Sciences, Salt Lake City, Kolkata-700106, India Contact author: priya@bose.res.in Nitesh Kumar Department of Condensed Matter and Materials Physics, S. N. Bose National Centre for Basic Sciences, Salt Lake City, Kolkata-700106, India Contact author: nitesh.kumar@bose.res.in
@@ 1 Introduction
Electrons have positive $m_{ij}^{*}$ while holes have negative $m_{ij}^{*}$. Since the second derivative of a function determines its curvature, $m_{ij}^{*}$ can be positive or negative depending on the curvature of the corresponding energy band crossing the Fermi energy ($E_{\mathrm{F}}$). DDCP can be realized if the integrated curvature of all the bands crossing $E_{\mathrm{F}}$ is positive in one direction and negative in the other. The FS, which provides a geometric representation of all points in reciprocal space where the band energy equals $E_{\mathrm{F}}$, encompasses all bands crossing $E_{\mathrm{F}}$. A single band material can exhibit DDCP if its FS has concave and convex characters along different directions [see Figure 1a]. This has been observed in $\mathrm{NaSn_{2}As_{2}}$,[1] which has a hyperboloid-shaped FS that has both positive and negative curvatures (dependent on direction). The DDCP arising from single carrier mechanism is termed “goniopolarity”, and the materials exhibiting it are called “goniopolar”. Another way to achieve DDCP is through multicarrier mechanism [see Figure 1(b-c)], where electrons and holes live in two distinct bands and dominate conduction along different crystallographic directions.[2, 6] This can be achieved when the FS exhibits highly anisotropic electron and hole pockets, such that the integrated curvature of the total FS has opposite signs along two different crystallographic directions.
@@ 2 Results and Discussion
CrSb crystallizes in a hexagonal NiAs-type structure with space group $P6_{3}/mmc$ and exhibits A-type antiferromagnetic order with a Néel temperature of approximately 703 K.[13, 14] The Cr moments are aligned parallel to the c-axis, coupled ferromagnetically in the ab-plane and antiferromagnetically along the c-axis [see Figure 2a]. The anisotropic arrangement of Sb atoms around the oppositely aligned Cr spins prevents the connection of opposite spin sublattices by simple lattice translations or inversion [see Figure 2b]. However, these sublattices can be connected either by a six-fold crystal rotation about the c-axis $(C_{\mathrm{6z}})$ combined with half the lattice translation along the c-axis $(\textit{{t}}_{1/2}=(0,0,c/2))$ or by a horizontal mirror $(M_{\mathrm{z}})$ passing through the Sb layer. These operations, when combined with time-reversal symmetry ($\tau$), yield the actual sublattice-transposing symmetries of the crystal: $C_{\mathrm{6z}}\textit{{t}}_{1/2}\tau$ and $M_{\mathrm{z}}\tau$. These symmetries give rise to four spin-degenerate nodal planes in the Brillouin zone (BZ) (see Supporting Information). Away from these high-symmetry nodal planes, the spin degeneracy is lifted, resulting in significant non-relativistic splitting of the energy bands (see Supporting Information).
@@
To examine the transport behavior of CrSb along different crystallographic directions, the coordinate axes have been defined on the hexagonal lattice of CrSb [see Figure 3a]. The x- and z-axes align with the crystallographic a- and c-directions whereas the y-axis is along the crystallographic $[01\bar{1}0]$ direction. Figure 3b shows the temperature dependence of in-plane $(\rho_{\mathrm{yy}})$ and cross-plane $(\rho_{\mathrm{zz}})$ longitudinal resistivity. The material shows metallic nature throughout the temperature range in both in-plane and cross-plane directions with almost coinciding resistivity values and similar residual resistivity ratios $(RRR\approx 3.7)$. The field dependent Hall resistivity data in different geometries are shown in Figure 3(c-e). When a magnetic field is applied within the hexagonal plane of the CrSb crystal, as illustrated in Figure 3c and 3d for $\textit{B}||\mathrm{x}$ and $\textit{B}||\mathrm{y}$ respectively, both the Hall resistivity $\rho_{\mathrm{zy}}$ and $\rho_{\mathrm{xz}}$ exhibit an almost linear dependence on the field at all temperatures with similar positive slopes. When the magnetic field is applied along the z-axis, the Hall resistivity $\rho_{\mathrm{yx}}$ deviates from linearity, especially at lower temperatures, and the slopes becomes predominantly negative. This non-linear behavior in Hall resistivity is often associated with the anomalous Hall effect [15] or a multicarrier transport mechanism.[16] Since the orientation of the Néel vector in CrSb precludes the occurrence of anomalous Hall effect,[17, 18, 19, 20] the non-linear behavior can be attributed to the multicarrier mechanism (see Supporting Information).[19, 21] Figure 3f presents the Hall coefficient, $R_{ijk}$, as a function of temperature for all Hall measurement geometries, evaluated at 0 T (low-field limit) and 9 T (high-field limit) using the equation: $R_{ijk}\left(B_{0}\right)=\left.\dfrac{d\rho_{ij}}{dB_{k}}\right|_{B_{0}}$. In a multicarrier system, the non-linear nature of Hall resistivity results in distinct low-field and high-field values for the Hall coefficient. The low-field value carries information of both carrier density and mobility, while the high-field value only carries the information of carrier denisty. [16, 22] Since both parameters play a crucial role in the overall conduction, the low-field value of the Hall coefficient is more reliable for determining the dominant conduction type in a material, with positive value indicating hole-dominant conduction and negative value indicating electron-dominant conduction. Across all temperatures, the low-field value of $R_{\mathrm{zyx}}$ and $R_{\mathrm{xzy}}$ remain positive, while that of $R_{\mathrm{yxz}}$ remains negative, inferring that the transport is dominated by electrons when the magnetic field is applied along the c-axis and by holes when the magnetic field is applied in the ab-plane of the hexagonal CrSb crystal. The sign of $R_{ijk}$ seems to be dictated only by the direction of the applied magnetic field, regardless of the direction of applied current. For two in-plane currents, $j||\mathrm{x}$ and $j||\mathrm{y}$, $R_{\mathrm{yxz}}$ and $R_{\mathrm{zyx}}$ have opposite signs, whereas for in-plane and out-of-plane currents, $j||\mathrm{y}$ and $j||\mathrm{z}$, $R_{\mathrm{zyx}}$ and $R_{\mathrm{xzy}}$ have the same sign. This happens because the trajectory of charge carrier (and hence the conduction direction) is determined by the direction of the magnetic field rather than the direction of the applied current in Hall effect measurements. In particular, the trajectory of charge carrier is given by the intersection of the FS by a plane perpendicular to the magnetic field as shown in Figure 1.[16, 22] The applied electric field (i.e; current) induces only a minute distortion in the FS and thus negligibly affects the trajectory of the charge carriers. Consequently, an in-plane magnetic field induces cross-plane conduction, whereas a cross-plane magnetic field induces in-plane conduction. This suggests that the in-plane conduction is electron-dominated whereas the cross-plane conduction is hole dominated. The identical sign and magnitude of $R_{\mathrm{zyx}}$ and $R_{\mathrm{xzy}}$ is a direct consequence of the magnetic point group $(6^{\prime}/m^{\prime}mm^{\prime})$ symmetry of CrSb which imposes the constraint $R_{\mathrm{zyx}}=R_{\mathrm{xzy}}$.[23, 19]
@@
$S_{ii}=-\frac{\pi^{2}K_{B}^{2}T}{3\left|e\right|}\left[\frac{1}{n(E)}\frac{dn(E)}{dE}+\frac{1}{\tau_{ii}(E)}\frac{d\tau_{ii}}{dE}\right.$
$\left.+m_{ii}^{\ast}\frac{d}{dE}\left(\frac{1}{m_{ii}^{\ast}}\right)\right]_{E=E_{\mathrm{F}}}.$ (2)
@@
We investigated the microscopic origin of direction-dependent changes in the nature of charge carriers involved in the transport of CrSb within DFT calculations. Figure 4a shows the band structure of CrSb, plotted along high-symmetry paths of the BZ [see the inset of Figure 4a] with spin-orbit coupling turned on. Since all these paths lie within the spin-degenerate nodal planes, protected by the sublattice-transposing symmetry of the CrSb crystal (see Supporting Information), the bands along these paths do not experience altermagnetic spin-splitting. Real crystals often deviate from perfect stoichiometry, which changes the position of the Fermi energy. This must, however, be considered in the DFT calculations to capture the trends seen in the experiment. We find that adding a small number of holes per unit cell ($\sim 0.13/\mathrm{u.c.}$) induces DDCP. In this case, the Fermi energy $E_{\mathrm{F}}$ is shifted below by 100 meV to $E^{\prime}_{\mathrm{F}}$, represented by the black dashed line in Figs. 4(a) and 4(b). To understand the unusual experimental results, we theoretically calculated the thermopower and the Hall coefficient tensor within the DFT framework using the Boltzmann transport equations and the relaxation-time approximation. The calculated Seebeck coefficient at 100 K is plotted as a function of the change in chemical potential in Figure 4b. It shows that direction-dependent conduction polarity exists within a narrow energy window ($\sim 14$ meV) around $E^{\prime}_{\mathrm{F}}$. The FS obtained at $E^{\prime}_{\mathrm{F}}$ [see Figure 5] is in good agreement with the FS observed in ARPES studies on CrSb,[25, 26, 27] confirming that Fermi energy is indeed shifted below its theoretical value in this system.
@@
The temperature-dependent in-plane ($S_{\mathrm{xx}}$) and cross-plane ($S_{\mathrm{zz}}$) Seebeck coefficients are presented in Figure 4c. Within the examined temperature range, the absolute value of $S_{\mathrm{xx}}$ exhibits a slight, monotonic increase with increasing temperature. Conversely, the absolute value of $S_{\mathrm{zz}}$ initially rises with temperature, reaches a broad peak, and then decreases. However, $S_{\mathrm{xx}}$ remains negative across this temperature range, while $S_{\mathrm{zz}}$ is positive, which is qualitatively consistent with the experimental observations. Figure 4d shows the computed Hall coefficient components as a function of temperature. While $R_{\mathrm{yxz}}$ remains negative throughout the chosen temperature range, $R_{\mathrm{yzx}}$ and $R_{\mathrm{zyx}}$ stay positive. The constraint $R_{\mathrm{yzx}}=R_{\mathrm{zyx}}$, dictated by the magnetic point group symmetry of CrSb, is also accurately reflected. The trends in Hall coefficients are also consistent with the experimental observations. These results suggest that the in-plane conduction is electron-dominated, while the cross-plane conduction is hole-dominated.
@@
To understand the mechanism of DDCP in CrSb, we calculated the FS at $E^{\prime}_{\mathrm{F}}$, as shown in Figure 5a. Three distinct Fermi pockets labeled $\alpha$, $\beta$, and $\gamma$ correspond to three different bands (#$\#1$, #$\#2$, and #$\#3$, respectively) that cross $E^{\prime}_{\mathrm{F}}$. Although band #$\#3$ does not cross $E^{\prime}_{\mathrm{F}}$ at $k_{z}=0$ plane, it crosses $E^{\prime}_{\mathrm{F}}$ at some finite value of $k_{z}$, as shown in Supporting Information Figure S7. The $\alpha$ and $\beta$ pockets appear to be nearly hyperboloid in shape, meaning they are closed in the $k_{\mathrm{x}}-k_{\mathrm{y}}$ plane and open along the $k_{\mathrm{z}}$ direction, apparently exhibiting both convex and concave features. Each of these pockets includes a small hemispherical sub-pocket around A point of the BZ. The $\gamma$ pocket consists of 12 symmetry-related similar sub-pockets. These sub-pockets are closed and have an anisotropic, bead-like shape, elongated along the $k_{\mathrm{z}}$-direction.
@@
Our DFT calculations predict that DDCP exists only within a small energy range ($\Delta E\approx 14\>\mathrm{meV}$) around $E^{\prime}_{\mathrm{F}}$. To test the validity of this, we doped CrSb with $2\%$ V and synthesized Cr0.98V0.02Sb. This introduces an additional 0.04 number of holes per unit cell which is expected to move the system out of the DDCP window. Figure 6a and 6b show the field-dependent Hall resistivity $\rho_{\mathrm{zy}}$ and $\rho_{\mathrm{yx}}$, respectively for Cr0.98V0.02Sb at temperatures 5 K and 300 K. While the hole doping has little effect on $\rho_{\mathrm{zy}}$, as evidenced by its similar positive slopes for CrSb and Cr0.98V0.02Sb (see 300 K data), it significantly modifies $\rho_{\mathrm{yx}}$, which retains non-linearity but now exhibits positive slopes, indicating hole-dominant conduction. This confirms that even a small amount of hole doping suppresses the DDCP and renders the system p-type, consistent with our DFT predictions.

### v2

@@
Banik Rai Department of Condensed Matter and Materials Physics, S. N. Bose National Centre for Basic Sciences, Salt Lake City, Kolkata-700106, India These authors contributed equally to this work. Krishnendu Patra Department of Condensed Matter and Materials Physics, S. N. Bose National Centre for Basic Sciences, Salt Lake City, Kolkata-700106, India These authors contributed equally to this work. Satyabrata Bera School of Physical Sciences, Indian Association for the Cultivation of Science, Jadavpur, Kolkata-700032, India Sk Kalimuddin School of Physical Sciences, Indian Association for the Cultivation of Science, Jadavpur, Kolkata-700032, India Kakan Deb Department of Condensed Matter and Materials Physics, S. N. Bose National Centre for Basic Sciences, Salt Lake City, Kolkata-700106, India Mintu Mondal School of Physical Sciences, Indian Association for the Cultivation of Science, Jadavpur, Kolkata-700032, India Priya Mahadevan Department of Condensed Matter and Materials Physics, S. N. Bose National Centre for Basic Sciences, Salt Lake City, Kolkata-700106, India Contact author: priya@bose.res.in Nitesh Kumar Department of Condensed Matter and Materials Physics, S. N. Bose National Centre for Basic Sciences, Salt Lake City, Kolkata-700106, India Contact author: nitesh.kumar@bose.res.in
@@ 1 Introduction
Electrons have positive $m_{ij}^{*}$ while holes have negative $m_{ij}^{*}$. Since the second derivative of a function determines its curvature, $m_{ij}^{*}$ can be positive or negative depending on the curvature of the corresponding energy band crossing the Fermi energy ($E_{\mathrm{F}}$). DDCP can be realized if the integrated curvature of all the bands crossing $E_{\mathrm{F}}$ is positive in one direction and negative in the other. The FS, which provides a geometric representation of all points in reciprocal space where the band energy equals $E_{\mathrm{F}}$, encompasses all bands crossing $E_{\mathrm{F}}$. A single band material can exhibit DDCP if its FS has concave and convex characters along different directions (see Figure 1a). This has been observed in $\mathrm{NaSn_{2}As_{2}}$,[1] which has a hyperboloid-shaped FS that has both positive and negative curvatures (dependent on direction). The DDCP arising from single carrier mechanism is termed “goniopolarity”, and the materials exhibiting it are called “goniopolar”. Another way to achieve DDCP is through multicarrier mechanism (see Figure 1(b-c)), where electrons and holes live in distinct bands and dominate conduction along different crystallographic directions.[2, 6] This can be achieved when the FS exhibits highly anisotropic electron and hole pockets, such that the integrated curvature of the total FS has opposite signs along two different crystallographic directions.
@@ 2 Results and Discussion
CrSb crystallizes in a hexagonal NiAs-type structure with space group $P6_{3}/mmc$ and exhibits A-type antiferromagnetic order with a Néel temperature of approximately 703 K.[13, 14] The Cr moments are aligned parallel to the c-axis, coupled ferromagnetically in the ab-plane and antiferromagnetically along the c-axis [see Figure 2a]. The anisotropic arrangement of Sb atoms around the oppositely aligned Cr spins prevents the connection of opposite spin sublattices by simple lattice translations or inversion (see Figure 2b). However, these sublattices can be connected either by a six-fold crystal rotation about the c-axis $(C_{\mathrm{6z}})$ combined with half the lattice translation along the c-axis $(\textit{{t}}_{1/2}=(0,0,c/2))$ or by a horizontal mirror $(M_{\mathrm{z}})$ passing through the Sb layer. These operations, when combined with time-reversal symmetry ($\tau$), yield the actual sublattice-transposing symmetries of the crystal: $C_{\mathrm{6z}}\textit{{t}}_{1/2}\tau$ and $M_{\mathrm{z}}\tau$. These symmetries give rise to four spin-degenerate nodal planes in the Brillouin zone (BZ) (see Figure S6a). Away from these high-symmetry nodal planes, the spin degeneracy is lifted, resulting in significant non-relativistic splitting of the energy bands (see Figure S6b).
@@
To examine the transport behavior of CrSb along different crystallographic directions, the coordinate axes have been defined on the hexagonal lattice of CrSb (see Figure 3a). The x- and z-axes align with the crystallographic a- and c-directions whereas the y-axis is along the crystallographic $[01\bar{1}0]$ direction. Figure 3b shows the temperature dependence of in-plane $(\rho_{\mathrm{yy}})$ and cross-plane $(\rho_{\mathrm{zz}})$ longitudinal resistivity. The material shows metallic nature throughout the temperature range in both in-plane and cross-plane directions with almost coinciding resistivity values and similar residual resistivity ratios $(RRR\approx 3.7)$. The field dependent Hall resistivity data in different geometries are shown in Figure 3(c-e). When a magnetic field is applied within the hexagonal plane of the CrSb crystal, as illustrated in Figure 3c and 3d for $\textit{B}||\mathrm{x}$ and $\textit{B}||\mathrm{y}$ respectively, both the Hall resistivity $\rho_{\mathrm{zy}}$ and $\rho_{\mathrm{xz}}$ exhibit an almost linear dependence on the field at all temperatures with similar positive slopes. When the magnetic field is applied along the z-axis, the Hall resistivity $\rho_{\mathrm{yx}}$ deviates from linearity, especially at lower temperatures, and the slopes becomes predominantly negative. This non-linear behavior in Hall resistivity is often associated with the anomalous Hall effect [15] or a multicarrier transport mechanism.[2] Since the orientation of the Néel vector in CrSb precludes the occurrence of anomalous Hall effect,[17, 18, 19, 20] the non-linear behavior can be attributed to the multicarrier mechanism (see Supporting Information).[19, 3] Figure 3f presents the Hall coefficient, $R_{ijk}$, as a function of temperature for all Hall measurement geometries, evaluated at 0 T (low-field limit) and 9 T (high-field limit) using the equation: $R_{ijk}\left(B_{0}\right)=\left.\dfrac{d\rho_{ij}}{dB_{k}}\right|_{B_{0}}$. In a multicarrier system, the non-linear nature of Hall resistivity results in distinct low-field and high-field values for the Hall coefficient. The low-field value carries information of both carrier density and mobility, while the high-field value only carries the information of carrier denisty. [2, 1] Since both parameters play a crucial role in the overall conduction, the low-field value of the Hall coefficient is more reliable for determining the dominant conduction type in a material, with positive value indicating hole-dominant conduction and negative value indicating electron-dominant conduction. Across all temperatures, the low-field value of $R_{\mathrm{zyx}}$ and $R_{\mathrm{xzy}}$ remain positive, while that of $R_{\mathrm{yxz}}$ remains negative, inferring that the transport is dominated by electrons when the magnetic field is applied along the c-axis and by holes when the magnetic field is applied in the ab-plane of the hexagonal CrSb crystal. The sign of $R_{ijk}$ is found to depend on whether the magnetic field is applied in-plane or out-of-plane rather than on the direction of the applied current. In particular, the trajectory of charge carrier is given by the intersection of the FS by a plane perpendicular to the magnetic field as shown in Figure 1.[2, 1] The applied electric field (i.e; current) induces only a minute distortion in the FS and thus negligibly affects the trajectory of the charge carriers. Consequently, an in-plane magnetic field induces cross-plane conduction, whereas a cross-plane magnetic field induces in-plane conduction. This suggests that the in-plane conduction is electron-dominated whereas the cross-plane conduction is hole dominated. The identical sign and magnitude of $R_{\mathrm{zyx}}$ and $R_{\mathrm{xzy}}$ is a direct consequence of the magnetic point group $(6^{\prime}/m^{\prime}mm^{\prime})$ symmetry of CrSb which imposes the constraint $R_{\mathrm{zyx}}=R_{\mathrm{xzy}}$.[23, 19]
$S_{ii}=-\frac{\pi^{2}K_{B}^{2}T}{3\left|e\right|}\left[\frac{1}{n(E)}\frac{dn(E)}{dE}+\frac{1}{\tau_{ii}(E)}\frac{d\tau_{ii}}{dE}+m_{ii}^{\ast}\frac{d}{dE}\left(\frac{1}{m_{ii}^{\ast}}\right)\right]_{E=E_{\mathrm{F}}}.$ (2)
@@
We investigated the microscopic origin of direction-dependent changes in the nature of charge carriers involved in the transport of CrSb within DFT calculations. Figure 4a shows the band structure of CrSb, plotted along high-symmetry paths of the BZ (see the inset of Figure 4a) with spin-orbit coupling (SOC) turned on. Since all these paths lie within the spin-degenerate nodal planes, which are protected by the sublattice-transposing symmetry of the CrSb crystal (see Supporting Information), the bands along these paths do not experience altermagnetic spin splitting in the absence of SOC.[25, 26, 27] However, when the SOC is included, the spin degeneracies in these nodal planes are lifted, leaving only a single nodal line along $\Gamma$–A. One finds a small spin polarization with in-plane spin components $\langle m_{x}\rangle$ and $\langle m_{y}\rangle$ along $\Gamma$–M, M–K, and K–$\Gamma$ directions (see Figure S7), a phenomenon referred to as weak altermagnetism.[27, 28] Real crystals often deviate from perfect stoichiometry, which changes the position of the Fermi energy. This must, however, be considered in the DFT calculations to capture the trends seen in the experiment. We find that adding a small number of holes per unit cell ($\sim 0.13/\mathrm{u.c.}$) induces DDCP. In this case, the Fermi energy $E_{\mathrm{F}}$ is shifted below by 100 meV to $E^{\prime}_{\mathrm{F}}$, represented by the black dashed line in Figure 4(a) and 4(b). To understand the unusual experimental results, we theoretically calculated the thermopower and the Hall coefficient tensor within the DFT framework using the Boltzmann transport equations and the relaxation-time approximation. The calculated Seebeck coefficient at 100 K is plotted as a function of the change in chemical potential in Figure 4b. It shows that the DDCP exists within a narrow energy window ($\sim 14$ meV) around $E^{\prime}_{\mathrm{F}}$. The FS obtained at $E^{\prime}_{\mathrm{F}}$ (see Figure 5) is in good agreement with the FS observed in ARPES studies on CrSb,[25, 26, 27] confirming that Fermi energy is indeed shifted below its theoretical value in this system.
@@
The temperature-dependent in-plane ($S_{\mathrm{xx}}$) and cross-plane ($S_{\mathrm{zz}}$) Seebeck coefficients are presented in Figure 4c. Within the examined temperature range, the absolute value of $S_{\mathrm{xx}}$ exhibits a slight, monotonic increase with increasing temperature. Conversely, the absolute value of $S_{\mathrm{zz}}$ initially rises with temperature, reaches a broad peak, and then decreases. However, $S_{\mathrm{xx}}$ remains negative across this temperature range, while $S_{\mathrm{zz}}$ is positive, which is qualitatively consistent with the experimental observations. We do not get quantitative agreement in the temperature dependence of the experimental data due to the approximations made while computing the Seebeck coefficients. The effects of phonons have been ignored in the description. [29, 30] These could enter through electron-phonon interactions as well as through phonon drag effects.[31] Figure 4d shows the computed Hall coefficient components as a function of temperature. While $R_{\mathrm{yxz}}$ remains negative throughout the chosen temperature range, $R_{\mathrm{xzy}}$ and $R_{\mathrm{zyx}}$ stay positive. The constraint $R_{\mathrm{xzy}}=R_{\mathrm{zyx}}$, dictated by the magnetic point group symmetry of CrSb, is also accurately reflected. The trends in Hall coefficients are also consistent with the experimental observations. These results suggest that the in-plane conduction is electron-dominated, while the cross-plane conduction is hole-dominated.
@@
To understand the mechanism of DDCP in CrSb, we calculated the FS at $E^{\prime}_{\mathrm{F}}$, as shown in Figure 5a. Three distinct Fermi pockets labeled $\alpha$, $\beta$, and $\gamma$ correspond to three different bands (#$\#1$, #$\#2$, and #$\#3$, respectively) that cross $E^{\prime}_{\mathrm{F}}$. Although band #$\#3$ does not cross $E^{\prime}_{\mathrm{F}}$ at $k_{z}=0$ plane, it crosses $E^{\prime}_{\mathrm{F}}$ at some finite value of $k_{z}$, as shown in Figure S8. The $\alpha$ and $\beta$ pockets appear to be nearly hyperboloid in shape, meaning they are closed in the $k_{\mathrm{x}}-k_{\mathrm{y}}$ plane and open along the $k_{\mathrm{z}}$ direction, apparently exhibiting both convex and concave features. Each of these pockets includes a small hemispherical sub-pocket around A point of the BZ. The $\gamma$ pocket consists of 12 symmetry-related similar sub-pockets. These sub-pockets are closed and have an anisotropic, bead-like shape, elongated along the $k_{\mathrm{z}}$-direction.
@@
Our DFT calculations predict that DDCP exists only within a small energy range ($\Delta E\approx 14\>\mathrm{meV}$) around $E^{\prime}_{\mathrm{F}}$. To test the validity of this, we doped CrSb with $2\%$ V and synthesized Cr0.98V0.02Sb. This introduces an additional 0.04 number of holes per unit cell which is expected to move the system out of the DDCP window. Figure 6a and 6b show the field-dependent Hall resistivity $\rho_{\mathrm{zy}}$ and $\rho_{\mathrm{yx}}$, respectively for Cr0.98V0.02Sb at temperatures 5 K and 300 K. While the hole doping has little effect on $\rho_{\mathrm{zy}}$, as evidenced by its similar positive slopes for CrSb and Cr0.98V0.02Sb (see 300 K data), it significantly modifies $\rho_{\mathrm{yx}}$, which retains non-linearity but now exhibits positive slopes, indicating hole-dominant conduction. A consistent result is also obtained in the thermopower measurement (see Figure S10) where the in-plane Seebeck coefficient, $S_{\mathrm{xx}}$, has changed sign and become positive for Cr0.98V0.02Sb. This confirms that even a small amount of hole doping suppresses the DDCP and renders the system p-type, consistent with our DFT predictions.